Question 1

Land rates on average cost K500 per year in 2010 and increase in value by 10% every year.

(i) What was the land rate in 2015?

(ii) If a man pays for a piece of land each year, how much will he have spent in total from 2010 to 2016?

(iii) In what year will the rate first exceed K1000?
1
Expert's answer
2019-08-15T10:03:39-0400

i)


"Q'=500(1+0.1)^5=K805.255"

ii)


"\frac{500((1+0.1)^7-1)}{0.1}=K4743.5855"

iii)


"500(1+0.1)^n>1000"

"(1+0.1)^n>2"

"n>\frac{\ln{2}}{\ln{1.1}}=7.3"

This year will be 2018.